OpenStudy (anonymous):

Solve the following system of equations by substitution: 4x + 3y = 12 y = x – 10 (Points : 4) (–4, 6) (6, –4) (–7, 3) (3, –7) HELPPPPPPPPPPPP PLEASEEEEEEEEEE

pooja195 (pooja195):

Since you already know what y equals plug it into E1 in place of y \[\huge~4x + 3(x-10) = 12 \]

pooja195 (pooja195):

Now solve for x ^

OpenStudy (anonymous):

WHATTTT??

pooja195 (pooja195):

distribute. 3(x−10)

OpenStudy (anonymous):

ohh so 3-10

OpenStudy (anonymous):

c maybe

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

?

pooja195 (pooja195):

3x-30 so 4x+3x-30=12 combine like terms.

OpenStudy (anonymous):

IMM SOO CONFUSED

pooja195 (pooja195):

http://www.mathsisfun.com/algebra/substitution.html read that

OpenStudy (anonymous):

ok I did

pooja195 (pooja195):

4x+3(x−10)=12

pooja195 (pooja195):

3x-30 so 4x+3x-30=12 combine like terms.